Consider events A and B are mutually exclusive, P(A)= 0.3 and P (B=0.6, p (A or B) is
1. 0.18
2. 0.09
3. 0
4. 0.9
5 0.3
1
Expert's answer
2020-11-02T18:32:32-0500

Solution


For two mutually exclusive events;


"P(A \cup B) = P(A) + P(B)""=0.6 +0.3 =0.9"

Ans: 0.9